OpenStudy (anonymous):

State the excluded value(s) of (x^2-5x+6)/(x^2-25) A. x cannot equal 5, –5 B. x cannot equal 2, 3 C. x cannot equal 5, –5, 2, 3 D. x cannot equal 5

jimthompson5910 (jim_thompson5910):

hint: the denominator cannot be zero

OpenStudy (ivettef365):

denominator = x^2 - 25, do you know how to solve this

OpenStudy (anonymous):

No I don't :S

jimthompson5910 (jim_thompson5910):

x^2-25 = 0 x = ?? or x = ??

OpenStudy (anonymous):

so the answer is b?

jimthompson5910 (jim_thompson5910):

x^2-25 = 0 (x-5)(x+5) = 0 x = ?? or x = ??

OpenStudy (anonymous):

a?

jimthompson5910 (jim_thompson5910):

correct